OpenStudy (anonymous):

Use the technique of completing the square to transform the equation into the form (x+c)^2=a 5x^2 + 20x -20=0

hero (hero):

1. Add 20 to both sides.

OpenStudy (anonymous):

Ok, I added 20 to both sides

hero (hero):

Post your result here so I can keep track.

hero (hero):

The reason why I'm asking you to post your steps here is so that if you make a mistake at any time, I can easily point it out for you.

OpenStudy (anonymous):

Ok, so when I added 20 to both sides I got 5x^2 + 20x=20. So then I think I divide by 5 So my new equation is x^2+ 4x=20?

hero (hero):

You divide BOTH sides by 5

OpenStudy (anonymous):

Sorry I knew that! So its x^2 +4x=4

hero (hero):

If you perform an operation on one side, you have to perform an operation on the other side.

hero (hero):

Yes, that is correct now

hero (hero):

Okay, so now what you have to do is add (b/2)^2 to both sides. In this case, b = 4, so add (4/2)^2 to both sides.

OpenStudy (anonymous):

Sorry im confused, math is hard for me

hero (hero):

Basically (4/2)^2 = (2)^2 = 4 So essentially I am asking you to add 4 to both sides.

hero (hero):

@cnh are you still here?

OpenStudy (anonymous):

Sorry, I am

OpenStudy (anonymous):

so x^2=24?

hero (hero):

How did you get from adding 4 to both sides, to that?

hero (hero):

You had: x^2 +4x=4 I asked you to add 4 to both sides.